if the graph of y=x²+mx+n passes through the points (1,12) and (3,28), what is the value of the produce mn?
Answer by ankor@dixie-net.com(22740)

if the graph of y=x²+mx+n passes through the points (1,12) and (3,28), what is the value of the product mn?
:
using the form ax^2 + bx + c = y,
a = 1; b = m; c = n
points
x = 1; y = 12
1^2 + 1m + n = 12
m + n = 12 - 1
m + n = 11
:
points:
x = 3; y = 28
3^2 + 3m + n = 28
9 + 3m + n = 28
3m + n = 28 - 9
3m + n = 19
:
Use elimination with these two equations
3m + n = 19
m + n = 11
-------------subtraction eliminates n, find m
2m = 8
m = 8/2
m = 4
:
find n using the equation: m + n = 11, m=4
n = 11 -4
n = 7
:
the equation for these points
y = x^2 + 4x + 7
:
"what is the value of the product mn?" Obviously it's 28